OpenStudy (anonymous):

Tom was asked how old he was. He replied "In 2 years I will be twice as old as I was five years ago.' How old is he ?

OpenStudy (mitu12):

he is 12yrs old solution: let x denote his age x+2=2(x-5) x+2=2x-10 2x-x=2+10 x=12

OpenStudy (anonymous):

so wait x=12?

OpenStudy (mitu12):

yes

OpenStudy (mitu12):

thank you

OpenStudy (anonymous):

no thank you

OpenStudy (mitu12):

no problem